# Mirage: Cross-Embodiment Zero-Shot Policy Transfer with Cross-Painting

### v1

@@ Abstract
The ability to reuse collected data and transfer trained policies between robots could alleviate the burden of additional data collection and training. While existing approaches such as pretraining plus finetuning and co-training show promise, they do not generalize to robots unseen in training. Focusing on common robot arms with similar workspaces and 2-jaw grippers, we investigate the feasibility of zero-shot transfer. Through simulation studies on 8 manipulation tasks, we find that state-based Cartesian control policies can successfully zero-shot transfer to a target robot after accounting for forward dynamics. To address robot visual disparities for vision-based policies, we introduce Mirage, which uses “cross-painting”—masking out the unseen target robot and inpainting the seen source robot—during execution in real time so that it appears to the policy as if the trained source robot were performing the task. Despite its simplicity, our extensive simulation and physical experiments provide strong evidence that Mirage can successfully zero-shot transfer between different robot arms and grippers with only minimal performance degradation on a variety of manipulation tasks such as picking, stacking, and assembly, significantly outperforming a generalist policy.
@@ I Introduction
Consider a scenario where substantial efforts are invested in training a vision-based policy on a Franka robot arm for a manipulation task. The conventional paradigm of policy transfer to another robot often involves collecting new data on the target robot and finetuning the pretrained policy or retraining with a combination of datasets. Collecting data and training policies for each task and physical robot embodiment is time-consuming and expensive. An exciting conjecture is that demonstration data from different robot arms can be combined to learn policies that are robust or more generalizable to different robots, such as Franka, Universal, ABB, KUKA, and Fanuc arms [19]. While scaling up shows great promise in in-distribution embodiments, transferring policies to unseen embodiments remains elusive.
@@
We present Mirage (Fig. LABEL:fig:teaser), a novel cross-embodiment policy transfer method that can zero-shot transfer policies trained on one source robot to an unseen target robot. Mirage decouples vision and control, allowing gaps between the source and target robot to be addressed separately. To address visual differences, Mirage employs “cross-painting” during execution, masking out the target robot and inpainting the source robot at the same pose, creating a mirage for the policy. Importantly, Mirage only requires knowledge of robot base coordinate frames for visual and state alignment. Mirage does not assume any demonstration data on the target robot or paired images or trajectories between the source and unseen target robots. To handle differences in control gains, Mirage pairs the source robot policy with a forward dynamics model and executes the action predicted by the policy on the target robot with a high-gain or blocking controller to accommodate varying control frequencies.
@@ II-A Transfer across embodiments
Transferring control dynamics. Previous work in control has investigated cross-robot transfer of dynamics models [18, 78, 39, 34] and alignment-based transfer [6, 64, 65] is a commonly used technique for trajectory tracking, where input-output data points for both robots are collected and then aligned using manifold alignment. The transformed data from the source robot can then be used as initialization to accelerate learning of the dynamics model of the target robot. Mirage does not focus on trajectory tracking. We fit a dynamics model to the source robot trajectory and use a high-gain or blocking controller on the target robot during execution.
Transfer learning and cross-domain imitation methods use finetuning to accelerate the learning process by leveraging data from other robots [99, 42, 123, 83]. For example, assuming access to a policy of the source robot, one can use Reinforcement Learning (RL) to finetune the visual encoder [82, 98], the value function [54], or the policy [60, 59] on the target robot. When learning from videos of other other agents without access to actions [85, 114, 7, 96, 61, 106, 108, 85, 3, 104], methods such as explicit retargeting [75, 95], inverse RL [116, 12], and goal-conditioned RL on a learned latent space [121] have been explored. Assuming isometry between domains, Fickinger et al. [30] train RL in the target domain by using a trajectory in the source domain as a pseudo-reward. Many other works assume data for both source and target robots on a proxy task, learn correspondences if not already available [2, 52, 119, 79], and then use the learned latent space or paired trajectories as auxiliary rewards [56, 2, 36, 91] or for adversarial training [38, 32, 111]. Unlike these methods, Mirage does not assume access to target robot data on a proxy task or paired trajectories, and directly reuses the source robot policy.
@@
Yang et al. [110] study both the control and visual domain gaps when transferring across robots. They use wrist cameras to minimize the observation differences of the robot, learn a multiheaded policy with robot-specific heads that capture separate dynamics, and use contrastive learning to align robot trajectories. In contrast, Mirage works with third-person camera(s) and does not assume any target robot data to jointly train a multi-robot policy. Another closely related work to ours is Robot-Aware Visual Foresight [41]. The authors use the known camera matrix and robot model to mask the robot pixels, train a video-prediction model that only predicts the world pixels, and use visual foresight [31, 25] during execution. Similar to their method, we use robot URDFs to compute robot pixels, but we not only mask out the target robot but also inpaint the source robot. By doing so, we do not impose restrictions on the policy class the source robot can use, allowing Mirage to work with state-of-the-art source robot policies such as diffusion policy [17] and successfully transfer them across real robots beyond pushing tasks [41].
@@ II-B Learning from large datasets
Beyond targeted effort to transfer robot policies, recent work has also explored use of large and diverse data [21, 49, 58, 27, 88, 29, 26, 102] to train visual encoders [71, 105, 63] and policies that are generalizable to new objects, scenes, tasks, as well as embodiments [1, 46, 62, 93, 97, 94, 46, 48, 80, 76, 89, 90, 4, 11, 10, 15, 23].  Dasari et al. [20] use a large dataset of multiple robots to pretrain a visual dynamics model. Most recently, several works [77, 8, 19, 73] pretrain large transformers and show that co-training or finetuning the models outperform policies trained only on in-domain data. However, RPT [77] uses joint space actions, preventing it from zero-shot transfer. While RT-X [19] and Octo [73] use Cartesian space actions, they do not align action spaces or condition on the robot coordinate frames and kinematics, preventing them from working zero-shot on a new robot setup as they do not know what action space to use. In this work, we conduct experiments that show that leveraging the robot URDFs and aligning the action spaces could enable even a single-robot specialist policy to generalize zero-shot to a different robot.
@@ III Problem Statement
We assume two robot arms, source $\mathcal{S}$ and target $\mathcal{T}$, with parallel-jaw grippers, known URDFs, and known isomorphic kinematics. We assume the manipulation tasks of interest are in both robots’ workspace and can be performed by both robots using similar strategies without collision. Our goal is to transfer a trained policy from the source robot to the target robot.
@@
We model each robot manipulation task of interest as a Markov Decision Process. We consider the setting where there is a policy $\pi_{\mathcal{S}}$ trained on a dataset of the source robot $D=\{(s_{1}^{\mathcal{S}},o_{1}^{\mathcal{S}},a_{1}^{\mathcal{S}},...,s_{H_{i}}^{\mathcal{S}},o_{H_{i}}^{\mathcal{S}})_{i}^{N}\}$ consisting of $N$ trajectories, where $s_{t}^{\mathcal{S}}$ is the proprioceptive state of the robot at timestep $t$, $o_{t}^{\mathcal{S}}$ is the third-person camera observation(s), and $a_{t}^{\mathcal{S}}$ is the action. Given a source policy $a_{t+1}^{\mathcal{S}}=\pi_{\mathcal{S}}(s_{t}^{\mathcal{S}},o_{t}^{\mathcal{S}})$, we would like to transform it into a policy $a_{t+1}^{\mathcal{T}}=\pi_{\mathcal{T}}(s_{t}^{\mathcal{T}},o_{t}^{\mathcal{T}})$ that takes as inputs the states and observations of the target robot without demonstration or finetuning trajectory data on the target robot.
Figure 0: Simulation Tasks and Robots. The simulation evaluation utilizes the Robosuite simulator with Lift, Stack, Can Pick-and-Place, Two Piece Assembly, and Square Peg Insertion tasks. Additionally, to study other policy classes, we evaluate policy transfers in ORBIT with a block lifting task, and in RLBench with 2 tasks: Lifting a lid, and Pushing a button to turn on a lamp. For all policies, the source robot is the Franka robot as shown in the first row, while the target robots for each of the tasks are shown in the second row.
@@ IV State-Based Transfer Experiments
UR5e IIWA Kinova Gen3 Sawyer Jaco UR5e IIWA Kinova Gen3 Sawyer $*{}^{*}$
@@
TABLE I: State-Based Policy Transfer Experiment Results. We evaluate state-based policies trained for each task using a Franka robot across five different robots equipped either with the original Franka gripper or with each target robot’s default gripper. Results suggest that most unseen target robots can successfully perform the tasks using the source robot as its guide for where to move its gripper. $*{}^{*}$aco has a 3-jaw gripper, which explains its lower success rates.
@@ IV-A Implementation Details
For ORBIT, we use the Lift task. We train an RL policy using PPO [87] on the Franka robot and evaluate on the UR5. Similar to Robosuite, we use the absolute pose controller to servo the desired pose at each step during execution. For RLBench, we use Coarse-to-Fine Q-attention [45] to learn the key poses and use its end-effector-pose-via-planning controller to reach the desired pose in an open-loop fashion. The policy also takes in camera observations and the scene point cloud. We study 2 tasks: take the lid off a saucepan (“Unlid Pan”) and turn on a lamp (“Lamp On”), and evaluate on the UR5 and Sawyer.
@@ VI Vision-Based Policy Transfer Experiments
    Can Mirage successfully zero-shot transfer trained vision-based policies from one robot to another?
@@ VI-A Simulation Experiments
Oracle Naive 0-shot Mirage Oracle Naive 0-shot Mirage
@@
TABLE II: Mirage Results on Transferring Vision-Based Policies in Simulation. For each task and robot arm combination, the Oracle represents the performance of a vision-based policy assuming access to a ground truth rendering of the source robot given the state of the target robot, the Naive 0-shot method directly passes the visual observation of the target robot to the policy, and Mirage uses cross-painting to generate the visual inputs for the policy. For each method, the first number represents the success rate when the target robot uses the source robot (Franka) gripper and the second number corresponds to using the target robot’s default gripper (Robotiq gripper).
@@
We first study the effect of cross-painting on mitigating the visual gap between robots. We focus on closed-loop policies. Based on Table I, we choose UR5e and Kinova Gen3 as two representative robots that show high potential for reusing the source robot policies. We use the ground-truth forward dynamics and compare the success rates of the target robot using cross-inpainting with no inpainting and ground-truth oracle rendering.
@@
To create a sim-to-sim gap, we choose Gazebo [53] as our renderer. We manually position spotlights in the Gazebo environment to simulate the lighting in Robosuite for similar robot renderings. Similar to Sec. IV, we first train the source robot policies with behavior cloning on the provided demonstration data for each task (200 demos for Lift and Can from Robomimic [68], 1000 demos for the other tasks from MimicGen [69]), using the LSTM architecture with the ResNet-18 visual encoder [68]. The policies utilize 84x84 images, and Mirage operates at approximately 40 Hz to cross-paint the images.
@@
Table II shows the results. Similar to Table I, transferring between robots but with the same gripper has higher success rates than transferring to a different gripper in most cases. For UR5e, Mirage has a 8-22% drop in performance compared to querying the source policy on an oracle rendering. There is a similar gap of at most 25% between using an oracle and using Mirage on Kinova Gen3. In all cases, Mirage significantly outperforms the naive 0-shot performance without any visual gap mitigation. This suggests that cross-painting can effectively bridge the visual differences of the robots.
@@ VI-B Physical Experiments
Task (Source Gripper) Source (Franka) Different Gripper Different Robot (UR5)
Baseline Octo Baseline 0-shot Octo Mirage Baseline 0-shot Octo Mirage
Tiger Pick-and-Place (Robotiq) 90% 70% 60% 70% 90% 50% 10% 90%
Open Drawer (Franka) 80% 40% 0% 20% 70% 0% 0% 60%
Stack Cup (Franka) 80% 30% 10% 20% 60% 0% 0% 50%
Toaster (Robotiq) 60% 20% 20% 0% 40% 0% 0% 30%
TABLE III: Mirage Results on Transferring Vision-Based Policies in Real. We evaluate Mirage on 4 tasks in 2 settings: Different Gripper: Transferring policies between the Franka gripper and the Robotiq 2F-85 gripper on a Franka robot. Different Robot: Using the Franka with either gripper as the source robot, and the UR5 robot with the Robotiq gripper as the target robot. Baseline/Baseline 0-shot: Separate Diffusion Policy models [17] trained on the source robot data for each task and evaluated on the source robot or zero-shot on the target embodiments. Octo: Octo Base model [73] finetuned on the source robot data from all tasks together and evaluated on the source robot or zero-shot on the target embodiments. Mirage: Evaluation of zero-shot transfer to the target embodiments using Mirage with the source policy being the corresponding baseline Diffusion Policy models.
@@
We study policy transfer in 2 different settings. The first setting is transferring between grippers only on the Franka robot. The second setting involves robot transfer, where we use the Franka with either gripper as the source robot, and the UR5 with the Robotiq 2F-85 as the target robot. We use one ZED 2 camera positioned from the side for each robot.
@@
For each task, we first use an Oculus controller to collect between 200-400 human demonstration trajectories on the source robot using VR teleoperation at 15 Hz [51]. We then train a separate diffusion policy [17] for each task, which we use as the source policy that we seek to transfer. On the target UR5 robot, we place the camera(s) at similar poses, with up to 4 cm differences from those in the source setup. Similar to Sec. VI-A, we use Gazebo as our renderer. We use a per-dimension linear forward dynamics model and use the demonstration data to fit the regression coefficients.
We compare Mirage to two baselines: naive 0-shot transfer and Octo finetuning [73]. Octo is a recently released state-of-the-art generalist policy for robotic manipulation, pre-trained on 800k robot episodes from the Open X-Embodiment dataset [19], including trajectories collected on Franka and UR5. For both methods, we use the known mapping between the robot coordinate frames to convert the end-effector poses (e.g., there is a ${}^{\circ}$ offset between the Franka gripper and Robotiq gripper when being installed on the Franka robot). We also use the same fitted forward dynamics model to account for controller differences, and use a blocking controller on the UR5 to reach the desired poses.
In the naive 0-shot transfer baseline, we directly apply the source policy (i.e., the task-specific diffusion policy trained on the source robot only) to the target embodiment for each task, after accounting for the coordinate frame and control differences as mentioned above. For the Octo model, we take the pretrained Base model (93M parameters) and finetune all weights on the combined 4-task demonstration data of the source robot instead of finetuning for each task separately. For both diffusion policy and Octo model, the image observation is the third-person camera view and the proprioception input is the Cartesian position of the end effector and the gripper position. The action output is also the Cartesian position and gripper position. For the Octo model, we use language goal conditioning. We evaluate both baseline models on all 3 robot setups: Source (Franka robot + Franka/Robotiq gripper), Different Gripper (Franka robot + the other gripper), and Different Robot (UR5 + Robotiq gripper).

### v3

@@ Abstract
The ability to reuse collected data and transfer trained policies between robots could alleviate the burden of additional data collection and training. While existing approaches such as pretraining plus finetuning and co-training show promise, they do not generalize to robots unseen in training. Focusing on common robot arms with similar workspaces and 2-jaw grippers, we investigate the feasibility of zero-shot transfer. Through simulation studies on 8 manipulation tasks, we find that state-based Cartesian control policies can successfully zero-shot transfer to a target robot after accounting for forward dynamics. To address robot visual disparities for vision-based policies, we introduce Mirage, which uses “cross-painting”—masking out the unseen target robot and inpainting the seen source robot—during execution in real time so that it appears to the policy as if the trained source robot were performing the task. Mirage applies to both first-person and third-person camera views and policies that take in both states and images as inputs or only images as inputs. Despite its simplicity, our extensive simulation and physical experiments provide strong evidence that Mirage can successfully zero-shot transfer between different robot arms and grippers with only minimal performance degradation on a variety of manipulation tasks such as picking, stacking, and assembly, significantly outperforming a generalist policy.
@@ I Introduction
Consider a scenario where substantial efforts are invested in training a vision-based policy on a Franka robot arm for a manipulation task. The conventional paradigm of policy transfer to another robot often involves collecting new data on the target robot and finetuning the pretrained policy or retraining with a combination of datasets. Collecting data and training policies for each task and physical robot embodiment are time-consuming and expensive. An exciting conjecture is that demonstration data from different robot arms can be combined to learn policies that are robust or more generalizable to different robots, such as Franka, Universal, ABB, KUKA, and Fanuc arms [19]. While scaling up shows great promise in in-distribution embodiments, transferring policies to unseen embodiments remains elusive.
@@
We present Mirage (Fig. LABEL:fig:teaser), a novel cross-embodiment policy transfer method that can zero-shot transfer policies trained on one source robot to an unseen target robot. Mirage decouples vision and control, allowing gaps between the source and target robot to be addressed separately. To address visual differences, Mirage employs “cross-painting” during execution, masking out the target robot and inpainting the source robot at the same pose, creating a “mirage” for the policy. Importantly, Mirage only requires knowledge of robot base coordinate frames for visual and state alignment. Mirage does not assume any demonstration data on the target robot or paired images or trajectories between the source and unseen target robots. Moreover, Mirage applies to both first-person and third-person camera views, and policies that take in both states and images as inputs or only images as inputs. To handle differences in control gains, Mirage pairs the source robot policy with a forward dynamics model and executes the action predicted by the policy on the target robot with a high-gain or blocking controller to accommodate varying control frequencies.
@@ II-A Transfer across embodiments
Transferring control dynamics. Previous work in control has investigated cross-robot transfer of dynamics models [18, 77, 38, 34]. Alignment-based transfer [6, 63, 64] is a commonly used technique for trajectory tracking, where input-output data points for both robots are collected and then aligned using manifold alignment. The transformed data from the source robot can then be used as initialization to accelerate learning of the dynamics model of the target robot. Mirage does not focus on trajectory tracking. We fit a dynamics model to the source robot trajectory and use a high-gain or blocking controller on the target robot during execution.
Transfer learning and cross-domain imitation methods use finetuning to accelerate the learning process by leveraging data from other robots [97, 41, 120, 81]. For example, assuming access to a policy of the source robot, one can use Reinforcement Learning (RL) to finetune the visual encoder [80, 96], the value function [53], or the policy [58, 59] on the target robot. When learning from videos of other agents without access to actions [83, 112, 7, 94, 60, 104, 106, 83, 3, 102], methods such as explicit retargeting [74, 93], inverse RL [114, 12], and goal-conditioned RL on a learned latent space [118] have been explored. Assuming isometry between domains, Fickinger et al. [30] train RL in the target domain by using a trajectory in the source domain as a pseudo-reward. Many other works assume data for both source and target robots on a proxy task, learn correspondences if not already available [2, 51, 116, 78], and then use the learned latent space or paired trajectories as auxiliary rewards [55, 2, 36, 89] or for adversarial training [37, 32, 109]. Unlike these methods, Mirage does not assume access to target robot data on a proxy task or paired trajectories, and directly reuses the source robot policy.
@@
Yang et al. [108] study both the control and visual domain gaps when transferring across robots. They use wrist cameras to minimize the observation differences of the robot, learn a multiheaded policy with robot-specific heads that capture separate dynamics, and use contrastive learning to align robot trajectories. In contrast, Mirage works with both first-person and third-person cameras and does not assume any target robot data to jointly train a multi-robot policy. Another closely related work to ours is Robot-Aware Visual Foresight [40]. The authors use the known camera matrix and robot model to mask the robot pixels, train a video-prediction model that only predicts the world pixels, and use visual foresight [31, 25] during execution. Similar to their method, we use robot URDFs to compute robot pixels, but we not only mask out the target robot but also inpaint the source robot. By doing so, we do not impose restrictions on the policy class the source robot can use, allowing Mirage to work with state-of-the-art source robot policies such as diffusion policy [17] and successfully transfer them across real robots beyond pushing tasks [40].
@@ II-B Learning from large datasets
Beyond targeted effort to transfer robot policies, recent work has also explored use of large and diverse data [21, 48, 57, 27, 86, 29, 26, 100] to train visual encoders [70, 103, 62] and policies that are generalizable to new objects, scenes, tasks, as well as embodiments [1, 45, 61, 91, 95, 92, 47, 79, 75, 87, 88, 4, 11, 10, 15, 23].  Dasari et al. [20] use a large dataset of multiple robots to pretrain a visual dynamics model. Most recently, several works [76, 8, 19, 72] pretrain large transformers and show that co-training or finetuning the models outperform policies trained only on in-domain data. However, RPT [76] uses joint space actions, preventing it from zero-shot transfer. While RT-X [19] and Octo [72] use Cartesian space actions, they do not align action spaces or condition on the robot coordinate frames and kinematics, preventing them from working zero-shot on a new robot setup as they do not know what action space to use. In this work, we conduct experiments that show that leveraging the robot URDFs and aligning the action spaces could enable even a single-robot specialist policy to generalize zero-shot to a different robot.
@@ III Problem Statement
We assume two robot arms, source $\mathcal{S}$ and target $\mathcal{T}$, with parallel-jaw grippers, known URDFs, and known isomorphic kinematics. We assume the manipulation tasks of interest are in both robots’ workspaces and can be performed by both robots using similar strategies without collision. Our goal is to transfer a trained policy from the source robot to the target robot.
@@
We model each robot manipulation task of interest as a Markov Decision Process. We consider the setting where there is a policy $\pi_{\mathcal{S}}$ trained on a dataset of the source robot $D=\{(s_{1}^{\mathcal{S}},o_{1}^{\mathcal{S}},a_{1}^{\mathcal{S}},...,s_{H_{i}}^{\mathcal{S}},o_{H_{i}}^{\mathcal{S}})_{i}^{N}\}$ consisting of $N$ trajectories, where $s_{t}^{\mathcal{S}}$ is the proprioceptive state of the robot at timestep $t$, $o_{t}^{\mathcal{S}}$ is the first- or third-person camera observation(s), and $a_{t}^{\mathcal{S}}$ is the action. Given a source policy action $a_{t+1}^{\mathcal{S}}=\pi_{\mathcal{S}}(s_{t}^{\mathcal{S}},o_{t}^{\mathcal{S}})$, we would like to transform it into a target policy action $a_{t+1}^{\mathcal{T}}=\pi_{\mathcal{T}}(s_{t}^{\mathcal{T}},o_{t}^{\mathcal{T}})$ that takes as inputs the states and observations of the target robot without demonstration or finetuning trajectory data on the target robot.
Figure 0: Simulation Tasks and Robots. The simulation evaluation utilizes the Robosuite simulator with Lift, Stack, Can Pick-and-Place, Two Piece Assembly, and Square Peg Insertion tasks. Additionally, to study other policy classes, we evaluate policy transfers in ORBIT with a block lifting task, and in RLBench with 2 tasks: Lifting a lid, and Pushing a button to turn on a lamp. For all policies, the source robot is the Franka robot as shown in the top row, while the target robots for each of the tasks are shown in the bottom row.
@@ IV State-Based Transfer Experiments
UR5e IIWA Kinova Gen3 Sawyer Jaco UR5e IIWA Kinova Gen3 Sawyer Jaco
@@
TABLE I: State-Based Policy Transfer Experiment Results. We evaluate state-based policies trained for each task using a Franka robot across five different robots equipped either with the original Franka gripper or with each target robot’s default gripper. Results suggest that most unseen target robots can successfully perform the tasks using the source robot as its guide for where to move its gripper. Jaco has a 3-jaw gripper, which explains its lower success rates.
@@ IV-A Implementation Details
For ORBIT, we use the Lift task. We train an RL policy using PPO [85] on the Franka robot and evaluate on the UR5. Similar to Robosuite, we use the absolute pose controller to servo to the desired pose at each step during execution. For RLBench, we use Coarse-to-Fine Q-attention [44] to learn the key poses and use its end-effector-pose-via-planning controller to reach the desired pose in an open-loop fashion. The policy also takes in camera observations and the scene point cloud. We study 2 tasks: take the lid off a saucepan (“Unlid Pan”) and turn on a lamp (“Lamp On”), and evaluate on the UR5 and Sawyer.
@@ VI Vision-Based Policy Transfer Experiments
    Can Mirage successfully zero-shot transfer trained vision-based policies from one robot to another? Does Mirage work with different camera views and policies?
@@ VI-A Simulation Experiments
Oracle Mirage  – CP Mirage Oracle Mirage  – CP Mirage
@@
TABLE II: Mirage Results on Transferring State-And-Vision-Based Policies in Simulation. For each task and robot arm combination, the source BC-RNN policy uses both robot gripper poses and images as inputs. Oracle represents the performance of the policy assuming access to a ground truth rendering of the source robot given the state of the target robot. Mirage  – CP ablates cross-painting and directly passes the visual observation of the target robot to the policy. Mirage uses cross-painting to generate the visual inputs for the policy. For each method, the first number represents the success rate when the target robot uses the source robot (Franka) gripper and the second number corresponds to using the target robot’s default gripper (Robotiq gripper).
Oracle Mirage  – CP Mirage Oracle Mirage  – CP Mirage
Lift 99% 99% / 96% 85% / 27% 96% / 96% 100% / 99% 74% / 14% 88% / 96%
Stack 97% 92% / 93% 0% / 0% 80% / 48% 95% / 96% 7% / 0% 76% / 52%
Can 60% 66% / 51% 0% / 0% 32% / 24% 51% / 36% 2% / 0% 24% / 20%
Two Piece Assembly 90% 88% / 87% 33% / 17% 64% / 40% 95% / 85% 38% / 7% 76% / 60%
Square 77% 70% / 48% 44% / 3% 40% / 20% 63% / 27% 19% / 0% 44% / 20%
TABLE III: Mirage Results on Transferring Vision-Only Policies in Simulation. For each task and robot arm combination, the source BC-RNN policy uses only the images as inputs. “Oracle” assumes access to a ground truth rendering of the source robot given the state of the target robot. Mirage  – CP ablates cross-painting and directly passes the visual observation of the target robot to the policy. Mirage uses cross-painting to generate the visual inputs for the policy. For each method, the first number represents the success rate when the target robot uses the source robot (Franka) gripper and the second number corresponds to using the target robot’s default gripper (Robotiq gripper).
@@
We first study the effect of cross-painting on mitigating the visual gap between robots. We focus on closed-loop policies. Based on Table I, we choose UR5e and Kinova Gen3 as two representative robots that show high potential for reusing the source robot policies. The end-effector poses $p_{\mathrm{r}}^{\mathcal{S}}$ and $p_{\mathrm{r}}^{\mathcal{T}}$ are aligned in Robosuite by default. To study the importance of cross-painting, we apply Mirage to policies that take in both states and images as inputs, as well as policies that only take in images as inputs. We use the ground-truth forward dynamics and compare the success rates of the target robot using cross-painting (“Mirage”), with no cross-painting (“Mirage  – CP”), and with ground-truth oracle rendering (“Oracle”).
@@
To create a sim-to-sim gap, we choose Gazebo [52] as our renderer. We manually position spotlights in the Gazebo environment to simulate the lighting in Robosuite for similar robot renderings. Similar to Sec. IV, we first train the source robot policies with behavior cloning (BC-RNN) on the provided demonstration data for each task (200 demos for Lift and Can from Robomimic [67], 1000 demos for the other tasks from MimicGen [68]), using the LSTM architecture with the ResNet-18 visual encoder [67]. The policies utilize 84x84 images, and Mirage operates at approximately 40 Hz to cross-paint the images.
@@
Table II shows the results for policies that take in both states and images as observations, and Table III shows those for policies that take in images only. Similar to Table I, transferring between robots but with the same gripper has higher success rates than transferring to a different gripper in most cases. For both UR5e and Kinova Gen3, Mirage has a performance drop of less than 25% compared to querying the source policy on an oracle rendering for most tasks. In all cases, Mirage significantly outperforms the naive 0-shot performance without any visual gap mitigation. This suggests that cross-painting can effectively bridge the visual differences of the robots.
@@ VI-B Physical Experiments
Policies Tasks Tiger in Bowl (Robotiq) Open Drawer (Franka) Stack Cup (Franka) Toaster (Robotiq)
Source T Grip T Rob Source T Grip T Rob Source T Grip T Rob Source T Grip T Rob
Policies that use both states and third-person images as observations:
$\pi_{\mathcal{S}}$ (Diffusion Policy) 90% 20% 0% 80% 0% 0% 80% 0% 0% 60% 0% 0%
$\pi_{\mathcal{S}}$+ Cross-Painting (CP) - 50% 0% - 10% 0% - 20% 0% - 0% 0%
$\pi_{\mathcal{S}}$+ State Alignment (SA) - 60% 50% - 0% 0% - 10% 0% - 20% 0%
$\pi_{\mathcal{S}}$+ SA + CP (Mirage) - 90% 90% - 70% 60% - 60% 50% - 40% 30%
Octo 70% 0% 0% 40% 0% 0% 30% 0% 0% 20% 0% 0%
Octo + State Alignment (SA) - 70% 10% - 20% 0% - 20% 0% - 0% 0%
Policies that use only third-person images as observations:
$\pi_{\mathcal{S}}$ (Diffusion Policy) 90% 20% 0% 70% 0% 0% 70% 0% 0% 50% 0% 0%
$\pi_{\mathcal{S}}$+ CP (Mirage) - 90% 60% - 60% 60% - 60% 40% - 30% 20%
TABLE IV: Mirage Results on Transferring Vision-Based Policies in Real with a Third-Person Camera. We evaluate Mirage on 4 tasks in 2 settings: Target (Different) Gripper (“T Grip”): Transferring policies between the Franka gripper and the Robotiq 2F-85 gripper on a Franka robot. Target (Different) Robot (“T Rob”): Using the Franka with either gripper as the source robot (indicated in the parentheses in the first row), and the UR5 robot with the Robotiq gripper as the target robot. Baseline $\pi_{\mathcal{S}}$ (Diffusion Policy): Separate Diffusion Policy models [17] trained on the source robot data for each task and evaluated on the source robot or zero-shot on the target embodiments. Octo: Octo Base model [72] finetuned on the source robot data from all tasks together and evaluated on the source robot or zero-shot on the target embodiments. Mirage: Evaluation of zero-shot transfer to the target embodiments using Mirage with the source policy being the corresponding baseline Diffusion Policy models. For policies that take in both states and images as inputs, Mirage applies both state alignment (SA) and cross-painting (CP). For policies that take in only images as inputs, Mirage applies cross-painting only. We perform ablations to study the importance of state alignment and cross-painting.
Policies Tasks Tiger (Robotiq) Drawer (Franka)
Source T Grip Source T Grip
Wrist Camera Only:
$\pi_{\mathcal{S}}$ (Diffusion Policy) 90% 0% 70% 0%
$\pi_{\mathcal{S}}$+ Cross-Painting (CP) - 10% - 0%
$\pi_{\mathcal{S}}$+ State Alignment (SA) - 0% - 0%
$\pi_{\mathcal{S}}$+ SA + CP (Mirage) - 70% - 40%
Wrist Camera + Third-Person Camera:
$\pi_{\mathcal{S}}$ (Diffusion Policy) 90% 0% 90% 0%
$\pi_{\mathcal{S}}$+ Cross-Painting (CP) - 0% - 0%
$\pi_{\mathcal{S}}$+ State Alignment (SA) - 0% - 0%
$\pi_{\mathcal{S}}$+ SA + CP (Mirage) - 80% - 60%
TABLE V: Mirage Results on Transferring Policies in Real that Use a First-Person (Wrist) Camera. We evaluate Mirage for transferring between grippers on 2 tasks, with the source gripper indicated in the parentheses in the first row. Target (Different) Gripper (“T Grip”): Transferring policies between the Franka gripper and the Robotiq 2F-85 gripper on a Franka robot. We study policies that use only the wrist camera or both wrist and third-person cameras. We perform ablations to study the importance of state alignment (SA) and cross-painting (CP).
@@
We study policy transfer in 2 different settings. The first setting is transferring between grippers only on the Franka robot. The second setting involves robot transfer, where we use the Franka with either gripper as the source robot, and the UR5 with the Robotiq 2F-85 as the target robot. We use a ZED 2 camera positioned from the side for each robot, and a ZED Mini as the wrist camera for each gripper.
@@
For each task, we first use an Oculus controller to collect between 200-400 human demonstration trajectories on the source robot using VR teleoperation at 15 Hz [50]. We then train 2 separate diffusion policies [17] for each task, one that uses both states and images as the inputs, and one that uses only images as inputs. They serve as the source policies that we seek to transfer. On the target UR5 robot, we place the camera(s) at similar poses, with up to 4 cm differences from those in the source setup. Similar to Sec. VI-A, we use Gazebo as our renderer. We use a per-dimension linear forward dynamics model and use the demonstration data to fit the regression coefficients. We use a blocking controller on the UR5 to reach the desired poses.
We note that on the Franka robot, there is a 45 difference in rotation between the Franka gripper and Robotiq gripper in their installations. Additionally, there is a 90 rotational difference in the coordinate frames of the Robotiq gripper in the base frame of the UR5 robot compared to that of the Franka robot. To study the value of state alignment (SA) $p_{\mathrm{r},t+1}^{\mathcal{T}}=T_{\mathcal{S}}^{\mathcal{T}}p_{\mathrm{r},t+1}^{\mathcal{S}}$ and cross-painting (CP) $o^{\mathcal{T}\rightarrow\mathcal{S}}$, we conduct ablations of both modules from Mirage. Additionally, as a baseline, we also directly apply the source policy $\pi_{\mathcal{S}}$ (the task-specific diffusion policy trained on the source robot only) to the target embodiment for each task, without accounting for either coordinate frame or robot visual differences.
Figure 3: Mirage applied to first-person wrist camera images and third-person front camera images. For each view, the top row shows the actual observations of the target robot during the rollout, and the bottom row shows the cross-painted images generated by Mirage.
We also compare Mirage to a finetuned multi-task policy, Octo [72]. Octo is a recently released state-of-the-art generalist policy for robotic manipulation, pre-trained on 800k robot episodes from the Open X-Embodiment dataset [19], including trajectories collected on Franka and UR5. We take the pretrained Base model (93M parameters) and finetune all weights on the combined 4-task demonstration data of the source robot with language goal conditioning. It uses a third-person camera image, and the proprioception input is the Cartesian position of the end effector and the gripper position.
Table IV shows the results for each task on all 3 robot setups: Source (Franka robot + Franka/Robotiq gripper), Different Gripper (Franka robot + the other gripper), and Different Robot (UR5 + Robotiq gripper). We can see that, for both gripper transfer and robot (and gripper) transfer, Mirage achieves strong zero-shot performance, significantly outperforming all baselines. Specifically, for policies that use both states and third-person images as inputs, state alignment and cross-painting are both crucial for task success on the target embodiments. This is not surprising since both transformations are needed to bring the test-time input back to the distribution of the training data. Comparing the two base models with state alignment only, task-specific diffusion policies achieve stronger performance on the source robot it is trained on, while Octo achieves better success rates on most tasks when the grippers are swapped, suggesting slightly more robustness to the visual changes. Still, Octo tends to miss more grasps or gets stuck, resulting in a lower success rate than Mirage by 20% to 50%. When being evaluated on the UR5, no baselines are able to achieve any success on the 3 more difficult tasks. For policies that take in only images as inputs, state alignment is unnecessary, and cross-painting alone can achieve strong results on the target embodiment, with only minor performance degradation compared to the source robot.